OpenStudy (anonymous):

my question say --> ' find the slope of the line containing the following two points (7/8, 3) & (-1/4, 1/4)

OpenStudy (anonymous):

do you need the fraction or decimal?

OpenStudy (anonymous):

1/4 - (-3) ? ____________ = _______ -1/4 - 7/8 ? 1/4 + 3 --------- -2/8 - 7/8 13/4 ----- -9/8 13/4 x (-8/9) = - 26/9

OpenStudy (anonymous):

that wasnt right ??? im tryna find the slope of the line of those two points

OpenStudy (anonymous):

can you give me all the questions?

OpenStudy (anonymous):

its says ' find the slope of the line containing the following two points ( 7/8 , 3 ) & ( -1/4, 1/4 )